OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job? A. 1,248 tiles B. 1,247 tiles C. 1,224 tiles D. 1,295 tiles

OpenStudy (anonymous):

Tiles required = area of room/area of tile => 109*103/3*3 = 1247.44 which is approximately equal to 1248

OpenStudy (anonymous):

Total Area of Room = ?? ZhangYan better explain rather than throwing the answer

OpenStudy (anonymous):

no 1248 is wrong.... 37*35=1295 (D)

OpenStudy (albert0898):

Hello! Thank you for posting your question on here. Since they give you the 2 sides of a rectangle and clearly ask for the measurement's inside, you should find the area. Side 1 (109) * Side 2 (103) = 11227 Each tile is 3 inches. Find the area of a square which is, side squared OR side * side. The area of the square tile is 9. To find how many tiles are needed, do 11227 divided by 9 and you will get 1247.444444 This number rounded is precisely equaled to 1,247; which is [B]

OpenStudy (anonymous):

Thx i just knew it wasnt 1248 :)

OpenStudy (anonymous):

thx for explaining it to me :)

OpenStudy (albert0898):

You're very welcome. Become my fan and receive instant help!